### v1

@@ 3.2 Worst-case Total Energy
$\overline{E}_{T}(x_{0},x_{tg},t_{f})\leq\frac{1}{t_{f}}\left\|B_{c}^{\dagger}(v-\tilde{x})\right\|_{2}^{2}+t_{f}\left(\sum_{i=1}^{p}\lambda_{i}\|V\|_{1}^{2}+p\right)+2\left\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right\|_{1},$ (27)
where $p$ is the dimension of the uncontrolled input $u_{uc}$, and $B_{uc}^{T}B_{uc}$ has the spectral decomposition $B_{uc}^{T}B_{uc}=V\Lambda V^{T}$, with $\Lambda=\operatorname*{diag}\{\lambda_{1},\ldots,\lambda_{p}\}$.
@@
$\overline{E}_{T}(x_{0},x_{tg},t_{f})\leq$
$\frac{1}{t_{f}}\left\|B_{c}^{\dagger}(v-\tilde{x})\right\|_{2}^{2}+\underbrace{\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}t_{f}\overline{u}_{uc}^{T}B_{uc}^{T}B_{uc}\overline{u}_{uc}}_{=T_{1}}+\underbrace{\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}2(\tilde{x}-v)^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}}_{=T_{2}}$
@@
where $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ implies the constraint $\|\overline{u}_{uc}\|_{\infty}\leq 1$ holds in the terms $T_{1}$ and $T_{2}$. We consider each term in (28) separately, beginning with $T_{1}$. Let the spectral decomposition of $B_{uc}^{T}B_{uc}$ be $B_{uc}^{T}B_{uc}=V\Lambda V^{T}$, where $\Lambda=\operatorname*{diag}\{\lambda_{1},\ldots,\lambda_{p}\}$ and $V$ is the orthonormal matrix of eigenvectors of $B_{uc}^{T}B_{uc}$. Next, define $q\coloneqq V^{T}\overline{u}_{uc}$. Using the sub-multiplicative property of norms, $\|q\|_{\infty}\leq\|V^{T}\|_{\infty}\|\overline{u}_{uc}\|_{\infty}\leq\|V\|_{1}$ since $\|\overline{u}_{uc}\|_{\infty}\leq 1$. Due to the use of the sub-multiplicative property, the set $\left\{\overline{u}_{uc}:\|\overline{u}_{uc}\|_{\infty}\leq 1\right\}$ is a subset of the set $\left\{q:\|q\|_{\infty}\leq\|V\|_{1}\right\}$, where $q=V^{T}\overline{u}_{uc}$. This property results in the following upper bound for $T_{1}$:
$T_{1}\leq t_{f}\sup_{\|q\|_{\infty}\leq\|V\|_{1}}q^{T}\Lambda q=t_{f}\sup_{\|q\|_{\infty}\leq\|V\|_{1}}\sum_{i=1}^{p}\lambda_{i}q_{i}^{2}=t_{f}\sum_{i=1}^{p}\lambda_{i}\|V\|_{1}^{2}.$ (29)
In (29), each component of the $q^{*}$ achieving the supremum takes on the maximum possible value, i.e., $q_{i}^{*}=\pm\|V\|_{1}$. Next, the term $T_{2}$ can be rewritten as the problem $\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}\beta^{T}\overline{u}_{uc}=\|\beta\|_{1}$, where $\beta=2B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(\tilde{x}-v)$ and the optimal $\overline{u}_{uc}$ for this particular term is $\overline{u}_{uc}^{*}=\operatorname*{sign}(\beta)=\operatorname*{sign}\left(2B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(\tilde{x}-v)\right)$. Then,
@@
Finally, $T_{3}$ is clearly maximized by each component of the uncontrolled input $u_{uc}$ taking on the maximal value of $+1$ or $-1$ at each time instant $t$, so that $u_{uc}^{T}(t)u_{uc}(t)=p$ for all $t$, and
@@
Proposition 2 thus provides an upper bound on the worst-case total energy, assuming the approximations for the nominal (20) and malfunctioning (26) energies are exact. In the next subsection, we consider the special case when control authority is lost over one actuator, i.e., when $p=1$. In this particular case, we derive a closed-form expression instead of an upper bound for $\overline{E}_{T}(x_{0},x_{tg},t_{f})$, assuming the above approximations are still exact. We subsequently use this closed-form expression to obtain a bound on the energetic resilience metric (12) in Section 4.
3.3 Loss of Authority over One Actuator
When control authority is lost over one actuator, we have $p=1$, $u_{uc}(t)\in\mathbb{R}$ and $B_{uc}\in\mathbb{R}^{n}$. Then, the worst-case total energy $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ is given by the following proposition.
@@ 3.3 Loss of Authority over One Actuator
$\overline{E}_{T}(x_{0},x_{tg},t_{f})=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}(v-\tilde{x})\right\|_{2}^{2}+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right)+2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right|.$ (32)
@@
$\overline{E}_{T}(x_{0},x_{tg},t_{f})=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}(v-\tilde{x})\right\|_{2}^{2}$ $+\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\biggl{\{}t_{f}B_{uc}^{T}B_{uc}\overline{u}_{uc}^{2}+2(v-\tilde{x})^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}$
$+\int_{0}^{t_{f}}u_{uc}^{2}(t)\mathrm{d}t\biggr{\}}$ (33)
The constraint $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ is equivalent to the constraints $-1\leq\overline{u}_{uc}\leq 1$ on $\overline{u}_{uc}$. Then, the second term inside the supremum reduces to
$\sup_{-1\leq\overline{u}_{uc}\leq 1}2(v-\tilde{x})^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}=2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right|,$
with the optimal $\overline{u}_{uc}^{*}$ for this particular term given by $\overline{u}_{uc}^{*}=\operatorname*{sign}\left(B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right)$. When $\operatorname*{sign}\left(B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right)\in\{-1,+1\}$, this choice of $\overline{u}_{uc}^{*}$ also maximizes the first term inside the supremum in (33), with $\sup_{-1\leq\overline{u}_{uc}\leq 1}t_{f}B_{uc}^{T}B_{uc}\overline{u}_{uc}^{2}=t_{f}\|B_{uc}\|_{2}^{2}$. Further, $\overline{u}_{uc}^{*}\in\{-1,+1\}$ implies that the uncontrolled input $u_{uc}^{*}(t)$ also takes on a constant value of $-1$ or $+1$ at each instant $t\in[0,t_{f}]$. This value for $u_{uc}^{*}$ maximizes the third term inside the supremum in (33), with $\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\int_{0}^{t_{f}}u_{uc}^{2}(t)\mathrm{d}t=t_{f}$.
On the other hand, if $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})=0$, the second term inside the supremum in (33) vanishes, leading to a smaller value for $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ than if this term did not vanish. The worst-case total energy is thus achieved only when $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\neq 0$. The uncontrolled input $u_{uc}^{*}$ maximizing each of the three terms inside the supremum in (33) is the same, and is a constant equal to $-1$ or $+1$ at almost every instant $t\in[0,t_{f}]$. Combining the results of each term above, we obtain the result (32). ∎
@@
We remind the reader that the primary results of this section rely on approximations made on the basis of Lemma 1. Further, the expressions for the different energies in (20), (26) and (27) depend on $v\equiv v(t_{f},u)$, with $\|v\|_{\infty}\leq\overline{v}$. These expressions play a key role in determining bounds on the energetic resilience metric (12), as shown in the next section.
4 Bounding the Energetic Resilience
In this section, we quantify the maximal additional energy used by the malfunctioning system (5), i.e., addressing Problem 2, for the particular case of loss of control authority over one actuator. Namely, we derive a bound on the energetic resilience metric (12), assuming the approximation for the nominal (20) and worst-case total (32) energies are exact. We note that the more general case where authority is lost over $p>1$ actuators can be addressed using (27). However, this bound is more conservative as (27) provides only an upper bound on the approximate worst-case total energy.
@@ 4 Bounding the Energetic Resilience
where we note that the norm inequality $\|v\|_{2}\leq\sqrt{n}\|v\|_{\infty}$ holds for $v\in\mathbb{R}^{n}$ [31, Chapter 5]. We then replace the constraint $\|\tilde{x}\|_{2}\leq R$ with $\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}$. Due to the norm inequality above, the set $\{d:\|\tilde{x}\|_{2}\leq R\}$ is a subset of the set $\{d:\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}$ where $\|v\|_{\infty}\leq\overline{v}$. This property results in the following upper bound on $r_{A}(t_{f},R)$:
@@
We can then use (34) to obtain the following bound on the resilience metric when $p=1$.
@@
$r_{A}(t_{f},R)\leq\frac{(R+\overline{v}\sqrt{n})^{2}}{t_{f}}\lambda_{\max}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)$ $+2(R+\overline{v}\sqrt{n})\left\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\right\|_{2}$
$+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right).$ (35)
@@
$r_{A}(t_{f},R)\leq t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right)$ $+2(R+\overline{v}\sqrt{n})\left\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\right\|_{2}$
$+\sup_{\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}}\frac{1}{t_{f}}(v-\tilde{x})^{T}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)(v-\tilde{x}),$
where the second term above follows from applying the Cauchy-Schwarz inequality to bound $2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right|$. Since $B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}$ is symmetric, we can use the Rayleigh inequality [31] and the constraint $\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}$ to bound the third term as
$\sup_{\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}}\frac{1}{t_{f}}(v-\tilde{x})^{T}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)(v-\tilde{x})\leq\frac{(R+\overline{v}\sqrt{n})^{2}}{t_{f}}\lambda_{\max}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right),$
@@
The bound in (35) thus quantifies the maximal additional energy needed for a nonlinear system which loses authority over one actuator, where we remind the reader of the approximations for the nominal (20) and worst-case total (32) energies. In the following section, we present an example that demonstrates the use of this metric in characterizing the increased energy required without significant conservatism.
5 Simulation Example

### v3

@@ 4.2. Worst-case Total Energy
$\overline{E}_{T}(x_{0},x_{tg},t_{f})\leq\frac{1}{t_{f}}\left\|B_{c}^{\dagger}(v-\tilde{x})\right\|_{2}^{2}+t_{f}\left(\sum_{i=1}^{p}{}_{i}\|V\|_{1}^{2}+p\right)+2\left\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right\|_{1},$ (39)
[…]
@@
$\overline{E}_{T}(x_{0},x_{tg},t_{f})$ $\leq\frac{1}{t_{f}}\left\|B_{c}^{\dagger}(v-\tilde{x})\right\|_{2}^{2}+\underbrace{\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}t_{f}\overline{u}_{uc}^{T}B_{uc}^{T}B_{uc}\overline{u}_{uc}}_{=T_{1}}+\underbrace{\sup_{\|\overline{u}_{uc}\|_{\infty}\leq 1}2(\tilde{x}-v)^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}}_{=T_{2}}$
@@
[…]
@@
Finally, $T_{3}$ is clearly maximized by each component of the uncontrolled input $u_{uc}$ taking on the maximal value of +1+1 or 1-1 at each time instant $t$, so that $u_{uc}^{T}(t)u_{uc}(t)=p$ for all $t$, and
@@
Proposition 3 thus provides an upper bound on the worst-case total energy, and we emphasize that this upper bound assumes the approximations for the nominal (32) and malfunctioning (38) energies are exact. In the next subsection, we consider the special case when control authority is lost over one actuator, i.e., when $p=1$. In this particular case, we derive a closed-form expression instead of an upper bound for $\overline{E}_{T}(x_{0},x_{tg},t_{f})$, assuming the above approximations are still exact. We subsequently use this closed-form expression to obtain a bound on the energetic resilience metric (12) in Section 5.
4.3. Loss of Authority over One Actuator
When control authority is lost over one actuator, we have $p=1$, $u_{uc}(t)\in\mdmathbb{R}$ and $B_{uc}\in\mdmathbb{R}^{n}$. Then, the worst-case total energy $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ is given by the following proposition.
@@ 4.3. Loss of Authority over One Actuator
[…]
@@
$\overline{E}_{T}(x_{0},x_{tg},t_{f})=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}(v-\tilde{x})\right\|_{2}^{2}+\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\biggl\{$ $t_{f}B_{uc}^{T}B_{uc}\overline{u}_{uc}^{2}+2(v-\tilde{x})^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}+\int_{0}^{t_{f}}u_{uc}^{2}(t)\mathrm{d}t\biggr\}$ (45)
The constraint $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ is equivalent to the constraints $1\leq\overline{u}_{uc}\leq 1$ on $\overline{u}_{uc}$. Then, the second term inside the supremum reduces to
[…]
On the other hand, if $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})=0$, the second term inside the supremum in (45) vanishes, leading to a smaller value for $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ than if this term did not vanish. The worst-case total energy is thus achieved only when $B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\neq 0$. The uncontrolled input $u_{uc}^{*}$ maximizing each of the three terms inside the supremum in (45) is the same, and is a constant equal to 1-1 or +1+1 at almost every instant $t\in[0,t_{f}]$. Combining the results of each term above, we obtain the result (44). ∎
@@
We remind the reader that the primary results of this section rely on approximations made on the basis of Lemma 1. Appendix B provides a classification of all non-approximate and approximate quantities derived in this paper. Further, the expressions for the different energies in (32), (38) and (39) depend on $v\equiv v(t_{f},u)$, with $\|v\|_{\infty}\leq\overline{v}$. These expressions play a key role in determining bounds on the energetic resilience metric (12), as shown in the next section.
5. Resilience of Nonlinear Systems
In this section, we quantify the maximal additional energy used by the malfunctioning system (5), i.e., addressing Problem 2, for the particular case of loss of control authority over one actuator. Namely, we derive a bound on the energetic resilience metric (12), assuming the approximation for the nominal (32) and worst-case total (44) energies are exact. We also discuss the reduction to the special case of linear driftless dynamics in (1) and recover an existing result from [27] in this case. We note that the more general case where authority is lost over $p>1$ actuators can be addressed using (39). However, this bound is more conservative as (39) provides only an upper bound on the approximate worst-case total energy.
@@ 5. Resilience of Nonlinear Systems
where we note that the norm inequality $\|v\|_{2}\leq\sqrt{n}\|v\|_{\infty}$ holds for $v\in\mdmathbb{R}^{n}$ [36, Chapter 5]. We then replace the constraint $\|\tilde{x}\|_{2}\leq R$ with $\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}$. Due to the norm inequality above, the set $\{d:\|\tilde{x}\|_{2}\leq R\}$ is a subset of the set $\{d:\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}\text{ where }\|v\|_{\infty}\leq\overline{v}$}. This property results in the following upper bound on $r_{A}(t_{f},R)$:
@@
We note that this analysis demonstrates why the metric used in [27] is unsuitable for nonlinear systems. Briefly, the metric in [27] uses a constraint of the form $\|\tilde{x}\|_{2}\geq R$ rather than $\|\tilde{x}\|_{2}\leq R$ in (12). As before, since (32) and (39) depend only on $v-\tilde{x}$, we would require a lower bound on $\|v-\tilde{x}\|_{2}$ to replace $\|\tilde{x}\|_{2}\geq R$. This lower bound can be derived as $\|v-\tilde{x}\|_{2}\geq R-\overline{v}\sqrt{n}$, but is meaningful only when $R>\overline{v}\sqrt{n}$. However, $\overline{v}$ in (28) can take on extremely large values, thus preventing $R>\overline{v}\sqrt{n}$ from being satisfied. Thus, we have introduced a different metric (12) in this paper for nonlinear systems.
We can then use (46) to obtain the following bound on the resilience metric.
@@
The energetic resilience $r_{A}(t_{f},R)$ is bounded from above as follows:
$r_{A}(t_{f},R)\leq t_{f}\Bigl(\sum_{i=1}^{p}{}_{i}\|V\|_{1}^{2}+p\Bigr)+\frac{(R+\overline{v}\sqrt{n})^{2}}{t_{f}}{}_{\max}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)+2n(R+\overline{v}\sqrt{n})\left\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\right\|_{2}.$ (47)
@@
$r_{A}(t_{f},R)$ $\leq\sup_{\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}}\overline{E}_{T}(x_{0},x_{tg},t_{f})-E_{N}^{*}(x_{0},x_{tg},t_{f})$
[…]
$\hskip 97.0001pt+\sup_{\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}}2\left\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}(v-\tilde{x})\right\|_{1}$
$\leq t_{f}\Bigl(\sum_{i=1}^{p}{}_{i}\|V\|_{1}^{2}+p\Bigr)+2n(R+\overline{v}\sqrt{n})\left\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\right\|_{2}$
[…]
where the second term above follows from the norm inequality $\|z\|_{1}\leq\sqrt{n}\|z\|_{2}$ for any vector $z\in\mdmathbb{R}^{n}$. Since $B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}$ is symmetric, we can use the Rayleigh inequality [36] and the constraint $\|v-\tilde{x}\|_{2}\leq R+\overline{v}\sqrt{n}$ to bound the third term as
[…]
@@
We also specialize this result to the case when $p=1$ in the following corollary.
Corollary 1.
[…]
Remark 2 (Reduction to Linear Driftless Systems).
We recall that setting $f(x(t))\equiv 0$ and $g(x(t))\equiv B$ recovers the driftless system (13) from (1). As discussed in Section 4, the quantity $v(t_{f},u)$ reduces to 0 so that $\overline{v}=0$. In this special case, equality is achieved in (31), so that (32) and (38) are no longer approximations, but hold with equality. Further, we can easily observe that the energies (32), (38), (39) and (44) reduce to the corresponding expressions (17), (18), (19) and (20) for the linear driftless system. The bound (48) on the energetic resilience metric also reduces to the corresponding bound (23) for linear driftless dynamics. We emphasize that each of these expressions for driftless dynamics hold without the need for approximations.
Finally, it is useful to investigate whether the bounds on the energetic resilience are tight. We now demonstrate that these bounds are achieved for the class of linear driftless systems with one state, losing authority over one input.
Proposition 6.
The bound (47) on the energetic resilience is exactly achieved for linear driftless systems of the form $\dot{x}(t)=b_{c}u_{c}(t)+b_{uc}u_{uc}(t)$, where all quantities are scalars.
Proof.
Note that in the special case of linear driftless systems, as per Remark 2, we recover all expressions derived in Section 3. Further, we know that $\overline{v}=0$ in this case. Let $B=[b_{c}~~b_{uc}]\in\mdmathbb{R}^{1\times 2}$. Then,
$E_{N}^{*}(x_{0},x_{tg},t_{f})=\frac{1}{t_{f}}\|B^{\dagger}\tilde{x}\|_{2}^{2}=\frac{\tilde{x}^{2}}{t_{f}}\|B^{\dagger}\|_{2}^{2}$
and
$\overline{E}_{T}(x_{0},x_{tg},t_{f})=\frac{\tilde{x}^{2}}{t_{f}b_{c}^{2}}+t_{f}(1+b_{uc}^{2})+2\left|\frac{b_{uc}\tilde{x}}{b_{c}^{2}}\right|.$
We note that these expressions are exact and not approximate, since the dynamics are linear and driftless. We also note that this value of $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ is achieved by the constant uncontrolled input
$u_{uc}^{*}(t)=\operatorname*{sign}\left(\frac{b_{uc}\tilde{x}}{b_{c}^{2}}\right)$
as shown in [27]. Substituting these into the definition of the metric,
$r_{A}(t_{f},R)$ $=t_{f}(1+b_{uc}^{2})+\sup_{|\tilde{x}|\leq R}\frac{\tilde{x}^{2}}{t_{f}}\left(\|B^{\dagger}\|_{2}^{2}-\frac{1}{b_{c}^{2}}\right)+2\left|\frac{b_{uc}\tilde{x}}{b_{c}^{2}}\right|$
$=t_{f}(1+b_{uc}^{2})+\frac{R^{2}}{t_{f}}\left(\|B^{\dagger}\|_{2}^{2}-\frac{1}{b_{c}^{2}}\right)+2R\left|\frac{b_{uc}}{b_{c}^{2}}\right|.$
It is clearly seen that this is simply the upper bound (34) with $\overline{v}=0$ since the system has linear driftless dynamics. The constant uncontrolled input above thus exactly achieves the upper bound in this case. ∎
@@
The bound in (48) thus quantifies the maximal additional energy needed for a nonlinear system which loses authority over one actuator, where we recall the approximations for the nominal (32) and worst-case total (44) energies. We also remind the reader that Appendix B provides a classification of non-approximate and approximate quantities derived in this paper. While the approximations may introduce an error in quantifying the actual energy used, we note that obtaining exact expressions for the energies (8) and (9), and subsequently the metric (12) is difficult due to the nonlinear nature of the system dynamics. We also note that a potential limitation of this metric is that the bound (48) may be conservative, and overestimate the maximal energy required. This may lead to conservative capacity design in practical systems, where resources such as fuel economy are important considerations. However, in the following section, we present three examples demonstrating the use of the energetic resilience metric in characterizing the increased energy required without significant conservatism, in the case of driftless, linear and nonlinear dynamical systems.
6. Simulation Examples